the area of a rectangular slab of sidewalk is 45 square feet. its length is 3 feet more than four times the width.find the length and the width.
Let width be x feet
Then length = 4x+3
Area = x*(4x+3)
which is equal to 45 (given)
Therefore
x*(4x+3)= 45
4x^2+3x = 45
4x^2 + 3x -45=0
4x^2 + 15x -12x -45=0
x*(4x+15) -3*(4x+15)=0
(x-3)(4x+15)=0
so x=3 or 15/4
Ans : Length = 15ft and width=3ft

Answer by mathstutor494(120)

Let width be x feet
Then length = 4x+3
Area = x*(4x+3)which is equal to 45 (given)
Therefore
x*(4x+3)= 45
4x^2+3x = 45
4x^2 + 3x -45=0
4x^2 + 15x -12x -45=0
x*(4x+15) -3*(4x+15)=0
(x-3)(4x+15)=0
so x=3 or 15/4
Ans : Length = 15ft and width=3ft